# Efficient Sign-Based Optimization: Accelerating Convergence via Variance Reduction

### v2

@@ 1 Introduction
where each $f_{i}(\cdot)$ is smooth and non-convex. This problem has been extensively studied in stochastic optimization (Johnson and Zhang, 2013; Defazio et al., 2014; Fang et al., 2018), but is less explored with sign-based methods. Previous literature proposes signSVRG (Chzhen and Schechtman, 2023) method to deal with the finite-sum problem, which achieves a convergence rate of $\mathcal{O}(m^{1/2}d^{1/2}T^{-1/2})$. However, its dependence on $m$ is sub-optimal, failing to match the $\mathcal{O}(m^{1/4}T^{-1/2})$ lower bound (Fang et al., 2018; Li et al., 2021) for problem (2). To address this gap, we propose the SSVR-FS algorithm, which periodically computes the exact gradient (Johnson and Zhang, 2013; Zhang et al., 2013) and incorporates it into the variance reduction estimator. In this way, we can achieve an improved convergence rate of $\mathcal{O}(m^{1/4}d^{1/2}T^{-1/2})$ for finite-sum problems.
@@ 2.2 Stochastic variance reduction methods
Stochastic variance reduction methods have gained significant attention in the optimization community in recent years. Among the pioneering approaches, the stochastic average gradient (SAG) method (Roux et al., 2012) and Stochastic Dual Coordinate Ascent (SDCA) algorithm (Shalev-Shwartz and Zhang, 2013) utilize a memory of previous gradients to ensure variance reduction, achieving linear convergence for strongly convex functions. To circumvent the need for storing gradients, the stochastic variance reduced gradient (SVRG) (Johnson and Zhang, 2013; Zhang et al., 2013) recalculates the full gradient periodically to enhance the accuracy of gradient estimators, maintaining linear convergence for strongly convex functions. Inspired by SAG and SVRG, Defazio et al. (2014) introduced the SAGA algorithm, which not only provides superior convergence rates but also supports proximal regularization. Subsequently, the stochastic recursive gradient algorithm (SARAH) (Nguyen et al., 2017) employs a simple recursive approach to update gradient estimators, ensuring better convergence for smooth convex functions.
For non-convex optimization, inspired by the SVRG algorithm, many methods (Reddi et al., 2016; Lei et al., 2017; Zhou et al., 2020) employ variance reduction techniques to design their algorithms and provide the corresponding convergence guarantees. More recent well-known advancements include the SPIDER (Fang et al., 2018) and SpiderBoost (Wang et al., 2019) methods, which improved the $\mathcal{O}(T^{-1/4})$ convergence rate of traditional SGD to $\mathcal{O}(T^{-1/3})$ under the average smoothness assumption. The convergence rate can be further improved to $\mathcal{O}(m^{1/4}T^{-1/2})$ for problems with a finite-sum structure, where $m$ represents the number of component functions. However, these methods typically require a huge batch size to ensure convergence. To avoid this limitation, the stochastic recursive momentum (STORM) method (Cutkosky and Orabona, 2019) introduces a momentum-based updating mechanism and an adaptive learning rate based on the stochastic gradients, achieving a convergence rate of $\mathcal{\tilde{O}}(T^{-1/3})$ without necessitating large batches. More recently, variance reduction techniques are widely employed in more complex problems to improve the existing convergence rates, such as compositional optimization (Wang et al., 2017a, b; Yuan et al., 2019; Zhang and Xiao, 2019; Jiang et al., 2022a, 2023), multi-level optimization (Chen et al., 2021; Zhang and Xiao, 2021; Jiang et al., 2022b, 2024b) and adaptive algorithms (Kavis et al., 2022; Jiang et al., 2024a).
@@ 3.2 Sign-based stochastic variance reduction for finite-sum structure
To handle the finite-sum problems, we retain the core structure of our SSVR method while incorporating elements from the SVRG (Johnson and Zhang, 2013; Zhang et al., 2013) approach. Specifically, we compute a full batch gradient at the first step and every $I$ iteration, i.e.,
@@ 4 Sign-based stochastic variance reduction with majority vote
This process is called majority vote (Bernstein et al., 2018), as each worker votes on the sign of the gradient, with the server tallying these votes and broadcasting the decision back to the nodes. However, the sign operation introduces bias in the estimation, and employing it twice can significantly amplify this bias, particularly in a heterogeneous environment. Previous analysis (Chen et al., 2020) indicates that signSGD fails to converge in the heterogeneous setting. To deal with this problem, we introduce an unbiased sign operation $\operatorname{S_{R}}(\cdot)$, which is defined below.
@@
For any vector $\mathbf{v}$ with $\|\mathbf{v}\|_{\infty}\leq R$, define the function mapping $\operatorname{S_{R}}(\mathbf{v})$ as:
$[\operatorname{S_{R}}(\mathbf{v})]_{k}=\begin{cases}+1,&\text{with probability }\frac{1}{2}+\frac{[\mathbf{v}]_{k}}{2R},\\ \\ -1,&\text{with probability }\frac{1}{2}-\frac{[\mathbf{v}]_{k}}{2R}.\end{cases}$ (5)
Remark: This operation provides an unbiased estimation of ${\mathbf{v}}/{R}$, such that $\mathbb{E}[\operatorname{S_{R}}(\mathbf{v})]={\mathbf{v}}/{R}$. It is worth noting that the function mapping is valid when $\|\mathbf{v}\|_{\infty}\leq R$, since the probability should always fall within $[0,1]$. For this purpose, we need to further assume that the gradient is bounded.
@@
4:         Draw sample $\xi_{t}^{j}$ and compute $\mathbf{v}_{t}^{j}=\nabla f_{j}(\mathbf{x}_{t};\xi_{t}^{j})+(1-\beta)\left(\mathbf{v}_{t-1}-\nabla f_{j}(\mathbf{x}_{t-1};\xi_{t}^{j})\right)$
5:         Option 1: Send $\operatorname{S_{R}}(\mathbf{v}_{t}^{j})$ to the parameter server, where $R=4G$
6:         Option 2: Send $\operatorname{S_{G}}(\hat{\mathbf{v}}_{t}^{j})$ to the parameter server, where $\hat{\mathbf{v}}_{t}^{j}=\Pi_{G}[\mathbf{v}_{t}^{j}]$
@@
8:         Option 1: Send $\mathbf{v}_{t}=\operatorname{Sign}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{R}}\left(\mathbf{v}_{t}^{j}\right)\right)$ to all nodes
9:         Option 2: Send $\mathbf{v}_{t}=\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{G}}\left(\hat{\mathbf{v}}_{t}^{j}\right)\right)$ to all nodes
@@
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\operatorname{Sign}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{R}}(\mathbf{v}_{t}^{j})\right).$
After applying $\operatorname{S_{R}}(\cdot)$, the output is a sign information, which can be transported between nodes efficiently. The complete algorithm, named SSVR with majority vote (SSVR-MV), is described in Algorithm 3 (with Option 1). Note that in Step 4, we set $\mathbf{v}_{1}^{j}=\nabla f(\mathbf{x}_{1};\xi_{1}^{j})$ when $t=1$. Next, we present the convergence guarantee for the proposed algorithm with the following assumption.
@@
To ensure each gradient estimator is bounded, we employ a projection operation $\hat{\mathbf{v}}_{t}^{j}=\Pi_{G}[\mathbf{v}_{t}^{j}]$, where $\Pi_{G}$ denotes the projection onto a ball of radius $G$. This allows us to utilize an unbiased sign mapping $\operatorname{S_{G}}(\hat{\mathbf{v}}_{t}^{j})$ before transmission to the parameter server. The revised algorithm is presented in Algorithm 3 (with Option 2), and the modifications lie in Steps 6 and 9. We now present the convergence guarantee for this modified approach below.
@@ Appendix C Proof of Theorem 3
$\begin{split}&\mathbb{E}\left[f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})\right]\\ \leq&2\eta R\sqrt{d}\mathbb{E}\left[\left\|\frac{\nabla f(\mathbf{x}_{t})}{R}-\frac{1}{n}\sum_{j=1}^{n}S_{R}(\mathbf{v}_{t}^{j})\right\|\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta R\sqrt{d}\mathbb{E}\left[\left\|\frac{\nabla f(\mathbf{x}_{t})}{R}-\frac{1}{nR}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\mathbb{E}\left[\left\|\frac{1}{n}\sum_{j=1}^{n}\left(S_{R}(\mathbf{v}_{t}^{j})-\frac{\mathbf{v}_{t}^{j}}{R}\right)\right\|\right]\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\sqrt{\mathbb{E}\left[\left\|\frac{1}{n}\sum_{j=1}^{n}\left(S_{R}(\mathbf{v}_{t}^{j})-\frac{\mathbf{v}_{t}^{j}}{R}\right)\right\|^{2}\right]}\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\sqrt{\frac{1}{n^{2}}\sum_{j=1}^{n}\mathbb{E}\left[\left\|\left(S_{R}(\mathbf{v}_{t}^{j})-\frac{\mathbf{v}_{t}^{j}}{R}\right)\right\|^{2}\right]}\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\sqrt{\frac{1}{n^{2}}\sum_{j=1}^{n}\mathbb{E}\left[\left\|S_{R}(\mathbf{v}_{t}^{j})\right\|^{2}\right]}\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+\frac{2\eta dR}{\sqrt{n}}-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2},\end{split}$ (12)
@@ Appendix D Proof of Theorem 4
$\begin{split}f(\mathbf{x}_{t+1})\leq&f(\mathbf{x}_{t})+\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L}{2}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|^{2}\\ \leq&f(\mathbf{x}_{t})-\eta\left\langle\nabla f(\mathbf{x}_{t}),\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}S_{G}(\hat{\mathbf{v}}_{t}^{j})\right)\right\rangle+\frac{\eta^{2}Ld}{2}\\ =&f(\mathbf{x}_{t})+\eta\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}S_{G}(\hat{\mathbf{v}}_{t}^{j})\right)\right\rangle-\eta\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{\eta^{2}Ld}{2}.\end{split}$
@@
$\begin{split}&\mathbb{E}\left[f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})\right]\\ \leq&\eta\mathbb{E}\left[\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}S_{G}(\hat{\mathbf{v}}_{t}^{j})\right)\right\rangle\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\eta\mathbb{E}\left[\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}S_{G}(\hat{\mathbf{v}}_{t}^{j})\right\rangle\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\eta\mathbb{E}\left[\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\hat{\mathbf{v}}_{t}^{j}\right\rangle\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\eta\mathbb{E}\left[\frac{1}{2}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{1}{2}\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\hat{\mathbf{v}}_{t}^{j}\right\|^{2}\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\frac{\eta}{2}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\hat{\mathbf{v}}_{t}^{j}\right\|^{2}\right]-\frac{\eta}{2}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\end{split}$ (14)

### v3

@@ 1 Introduction
where each $f_{i}(\cdot)$ is smooth and non-convex. This problem has been extensively studied in stochastic optimization (Zhang et al., 2013; Defazio et al., 2014; Fang et al., 2018), but is less explored with sign-based methods. Previous literature proposes signSVRG (Chzhen and Schechtman, 2023) method to deal with the finite-sum problem, which achieves a convergence rate of $\mathcal{O}(m^{1/2}d^{1/2}T^{-1/2})$. However, its dependence on $m$ is sub-optimal, failing to match the $\mathcal{O}(m^{1/4}T^{-1/2})$ lower bound (Fang et al., 2018; Li et al., 2021) for problem (2). To address this gap, we propose the SSVR-FS algorithm, which periodically computes the exact gradient (Zhang et al., 2013; Johnson and Zhang, 2013) and incorporates it into the variance reduction estimator. In this way, we can achieve an improved convergence rate of $\mathcal{O}(m^{1/4}d^{1/2}T^{-1/2})$ for finite-sum problems.
@@ 2.2 Stochastic variance reduction methods
Stochastic variance reduction methods have gained significant attention in the optimization community in recent years. Among the pioneering approaches, the stochastic average gradient (SAG) method (Roux et al., 2012) and Stochastic Dual Coordinate Ascent (SDCA) algorithm (Shalev-Shwartz and Zhang, 2013) utilize a memory of previous gradients to ensure variance reduction, achieving linear convergence for strongly convex functions. To circumvent the need for storing gradients, the stochastic variance reduced gradient (SVRG) (Zhang et al., 2013; Johnson and Zhang, 2013) recalculates the full gradient periodically to enhance the accuracy of gradient estimators, maintaining linear convergence for strongly convex functions. Inspired by SAG and SVRG, Defazio et al. (2014) introduced the SAGA algorithm, which not only provides superior convergence rates but also supports proximal regularization. Subsequently, the stochastic recursive gradient algorithm (SARAH) (Nguyen et al., 2017) employs a simple recursive approach to update gradient estimators, ensuring better convergence for smooth convex functions.
For non-convex optimization, inspired by the SVRG algorithm, many methods (Reddi et al., 2016; Lei et al., 2017; Zhou et al., 2020) employ variance reduction to design their algorithms and provide the corresponding convergence guarantees. More recent well-known advancements include the SPIDER (Fang et al., 2018) and SpiderBoost (Wang et al., 2019) methods, which improved the $\mathcal{O}(T^{-1/4})$ convergence rate of traditional SGD to $\mathcal{O}(T^{-1/3})$ under the average smoothness assumption. The convergence rate can be further improved to $\mathcal{O}(m^{1/4}T^{-1/2})$ for problems with a finite-sum structure, where $m$ represents the number of component functions. However, these methods typically require a huge batch size to ensure convergence. To avoid this limitation, the stochastic recursive momentum (STORM) method (Cutkosky and Orabona, 2019) introduces a momentum-based updating and an adaptive learning rate based on the stochastic gradients, achieving a convergence rate of $\mathcal{\tilde{O}}(T^{-1/3})$ without necessitating large batches. More recently, variance reduction techniques are widely employed in more complex problems to improve the existing convergence rates, such as compositional optimization (Wang et al., 2017a, b; Yuan et al., 2019; Jiang et al., 2022a, 2023), multi-level optimization (Chen et al., 2021; Zhang and Xiao, 2021; Jiang et al., 2022b, 2024b), adaptive algorithms (Kavis et al., 2022; Jiang et al., 2024a), and distributionally robust optimization (Yu et al., 2024).
@@ 3.2 Sign-based stochastic variance reduction for finite-sum structure
To handle the finite-sum problems, we retain the core structure of our SSVR method while incorporating elements from the SVRG (Zhang et al., 2013; Johnson and Zhang, 2013) approach. Specifically, we compute a full batch gradient at the first step and every $I$ iteration, i.e.,
@@ 4 Sign-based stochastic variance reduction with majority vote
This process is called majority vote (Bernstein et al., 2018), as each worker votes on the sign of the gradient, with the server tallying these votes and broadcasting the decision back to the nodes. However, the sign operation introduces bias in the estimation, and employing it twice can significantly amplify this bias, particularly in a heterogeneous environment. Previous analysis (Chen et al., 2020) indicates that signSGD fails to converge in the heterogeneous setting. To deal with this problem, we introduce an unbiased sign operation $\operatorname{S_{\textit{R}}}(\cdot)$, which is defined below.
@@
For any vector $\mathbf{v}$ with $\|\mathbf{v}\|_{\infty}\leq R$, define the function mapping $\operatorname{S_{\textit{R}}}(\mathbf{v})$ as:
$[\operatorname{S_{\textit{R}}}(\mathbf{v})]_{k}=\begin{cases}+1,&\text{with probability }\frac{1}{2}+\frac{[\mathbf{v}]_{k}}{2R},\\ \\ -1,&\text{with probability }\frac{1}{2}-\frac{[\mathbf{v}]_{k}}{2R}.\end{cases}$ (5)
Remark: This operation provides an unbiased estimation of ${\mathbf{v}}/{R}$, such that $\mathbb{E}[\operatorname{S_{\textit{R}}}(\mathbf{v})]={\mathbf{v}}/{R}$. It is worth noting that the function mapping is valid when $\|\mathbf{v}\|_{\infty}\leq R$, since the probability should always fall within $[0,1]$. For this purpose, we need to further assume that the gradient is bounded.
@@
4:         Draw sample $\xi_{t}^{j}$ and compute $\mathbf{v}_{t}^{j}=\nabla f_{j}(\mathbf{x}_{t};\xi_{t}^{j})+(1-\beta)\left(\mathbf{v}_{t-1}^{j}-\nabla f_{j}(\mathbf{x}_{t-1};\xi_{t}^{j})\right)$
5:         Option 1: Send $\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})$ to the parameter server, where $R=4G$
6:         Option 2: Send $\operatorname{S_{\textit{G}}}(\hat{\mathbf{v}}_{t}^{j})$ to the parameter server, where $\hat{\mathbf{v}}_{t}^{j}=\Pi_{G}[\mathbf{v}_{t}^{j}]$
@@
8:         Option 1: Send $\mathbf{v}_{t}=\operatorname{Sign}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{R}}}\left(\mathbf{v}_{t}^{j}\right)\right)$ to all nodes
9:         Option 2: Send $\mathbf{v}_{t}=\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{G}}}\left(\hat{\mathbf{v}}_{t}^{j}\right)\right)$ to all nodes
@@
$\mathbf{x}_{t+1}=\mathbf{x}_{t}-\eta\operatorname{Sign}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})\right).$
After applying $\operatorname{S_{\textit{R}}}(\cdot)$, the output is a sign information, which can be transported between nodes efficiently. The complete algorithm, named SSVR with majority vote (SSVR-MV), is described in Algorithm 3 (with Option 1). Note that in Step 4, we set $\mathbf{v}_{1}^{j}=\nabla f(\mathbf{x}_{1};\xi_{1}^{j})$ when $t=1$. Next, we present the convergence guarantee for the proposed algorithm with the following assumption.
@@
To ensure each gradient estimator is bounded, we employ a projection operation $\hat{\mathbf{v}}_{t}^{j}=\Pi_{G}[\mathbf{v}_{t}^{j}]$, where $\Pi_{G}$ denotes the projection onto a ball of radius $G$. This allows us to utilize an unbiased sign mapping $\operatorname{S_{\textit{G}}}(\hat{\mathbf{v}}_{t}^{j})$ before transmission to the parameter server. The revised algorithm is presented in Algorithm 3 (with Option 2), and the modifications lie in Steps 6 and 9. We now present the convergence guarantee for this modified approach below.
@@ Appendix C Proof of Theorem 3
$\begin{split}&\mathbb{E}\left[f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})\right]\\ \leq&2\eta R\sqrt{d}\mathbb{E}\left[\left\|\frac{\nabla f(\mathbf{x}_{t})}{R}-\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})\right\|\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta R\sqrt{d}\mathbb{E}\left[\left\|\frac{\nabla f(\mathbf{x}_{t})}{R}-\frac{1}{nR}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\mathbb{E}\left[\left\|\frac{1}{n}\sum_{j=1}^{n}\left(\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})-\frac{\mathbf{v}_{t}^{j}}{R}\right)\right\|\right]\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\sqrt{\mathbb{E}\left[\left\|\frac{1}{n}\sum_{j=1}^{n}\left(\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})-\frac{\mathbf{v}_{t}^{j}}{R}\right)\right\|^{2}\right]}\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\sqrt{\frac{1}{n^{2}}\sum_{j=1}^{n}\mathbb{E}\left[\left\|\left(\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})-\frac{\mathbf{v}_{t}^{j}}{R}\right)\right\|^{2}\right]}\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+2\eta R\sqrt{d}\sqrt{\frac{1}{n^{2}}\sum_{j=1}^{n}\mathbb{E}\left[\left\|\operatorname{S_{\textit{R}}}(\mathbf{v}_{t}^{j})\right\|^{2}\right]}\\ &\quad-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&2\eta\sqrt{d}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{t}^{j}\right\|\right]+\frac{2\eta dR}{\sqrt{n}}-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|_{1}\right]+\frac{\eta^{2}Ld}{2},\end{split}$ (12)
@@ Appendix D Proof of Theorem 4
$\begin{split}f(\mathbf{x}_{t+1})\leq&f(\mathbf{x}_{t})+\left\langle\nabla f(\mathbf{x}_{t}),\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\rangle+\frac{L}{2}\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\|^{2}\\ \leq&f(\mathbf{x}_{t})-\eta\left\langle\nabla f(\mathbf{x}_{t}),\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{G}}}(\hat{\mathbf{v}}_{t}^{j})\right)\right\rangle+\frac{\eta^{2}Ld}{2}\\ =&f(\mathbf{x}_{t})+\eta\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{G}}}(\hat{\mathbf{v}}_{t}^{j})\right)\right\rangle-\eta\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{\eta^{2}Ld}{2}.\end{split}$
@@
$\begin{split}&\mathbb{E}\left[f(\mathbf{x}_{t+1})-f(\mathbf{x}_{t})\right]\\ \leq&\eta\mathbb{E}\left[\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\operatorname{S_{1}}\left(\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{G}}}(\hat{\mathbf{v}}_{t}^{j})\right)\right\rangle\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\eta\mathbb{E}\left[\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\operatorname{S_{\textit{G}}}(\hat{\mathbf{v}}_{t}^{j})\right\rangle\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\eta\mathbb{E}\left[\left\langle\nabla f(\mathbf{x}_{t}),\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\hat{\mathbf{v}}_{t}^{j}\right\rangle\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\eta\mathbb{E}\left[\frac{1}{2}\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}+\frac{1}{2}\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\hat{\mathbf{v}}_{t}^{j}\right\|^{2}\right]-\eta\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\\ \leq&\frac{\eta}{2}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})-\frac{1}{n}\sum_{j=1}^{n}\hat{\mathbf{v}}_{t}^{j}\right\|^{2}\right]-\frac{\eta}{2}\mathbb{E}\left[\left\|\nabla f(\mathbf{x}_{t})\right\|^{2}\right]+\frac{\eta^{2}Ld}{2}\end{split}$ (14)